# q11 - DEBUG function futureValue(principal,N,interestRate)
# - This function currently returns INCORRECT answers
# - Find and fix the errors in the function below
#
# - the program accepts 3 parameters:
# --- principal - float (starting amount)
# --- N - integer (number of years)
# --- interestRate - float (value less than 1)
#
# - futureValue should calculate the future value of the principal
# when invested at the interestRate for N years. Interest is
# compounded at the end of each year.
#
# - Example:
# principal = 1000.00
# N = 3
# interestRate = .05 (this is 5%)
#
# The value at the end of each year is:
# N1: 1000.00 * 1.05 = 1050.00
# N2: 1050.00 * 1.05 = 1102.50
# N3: 1102.50 * 1.05 = 1157.625
#
# return 1157.625
Answer #1

In case of any query do comment. Please rate answer as well. Thanks

Code:

# q11 - DEBUG function futureValue(principal,N,interestRate)
# - This function currently returns INCORRECT answers
# - Find and fix the errors in the function below
#
# - the program accepts 3 parameters:
# --- principal - float (starting amount)
# --- N - integer (number of years)
# --- interestRate - float (value less than 1)
#
# - futureValue should calculate the future value of the principal
# when invested at the interestRate for N years. Interest is
# compounded at the end of each year.
#
# - Example:
# principal = 1000.00
# N = 3
# interestRate = .05 (this is 5%)
#
# The value at the end of each year is:
# N1: 1000.00 * 1.05 = 1050.00
# N2: 1050.00 * 1.05 = 1102.50
# N3: 1102.50 * 1.05 = 1157.625
#
# return 1157.625

def futureValue(principal,N,interestRate):
amount = principal
for i in range(N):
amount = amount + (amount * interestRate)
print("N{}: Amount: {}".format(i+1,amount))
return amount

#main driver function
amount = futureValue(1000,3,0.05)
print("Future Value: ", amount)

==========Screen shot of the code for indentation=======

WONO main.py 1 # 911 - DEBUG function futureValue(principal, n, interestRate) 2 # - This function currently returns INCORRECT

Output:

N1: Amount: 1050.0 N2: Amount: 1102.5 N3: Amount: 1157.625 Future Value: 1157.625 ... Program finished with exit code 0 Press